### v1

@@ 4.1 Block nullification
Since $\bm{\Omega}_{i}^{(\ell)}\in\mathbb{R}^{2k\times s}$, it has a nullspace of dimension at least $s-2k>k$. Therefore, we can obtain a matrix $\bm{P}_{i}^{(\ell)}$ whose columns form an orthonormal basis for the nullspace of $\bm{\Omega}_{i}^{(\ell)}$. Then, $\bm{\Omega}_{i}^{(\ell)}\bm{P}_{i}^{(\ell)}=\bm{0}$. Moreover, for all $j\neq i$, the unitary invariance of Gaussian matrices together with the mutual independence of $\bm{\Omega}_{i}$ and $\bm{\Omega}_{j}$ imply that $\bm{\Omega}_{j}^{(\ell)}\bm{P}_{i}^{(\ell)}$ are independent Gaussian matrices. Therefore, if we define
@@ 4.2 Low-rank approximation via projection-cost-preserving sketches
Define $\bm{B}_{1}=\bm{U}_{1}\bm{\Sigma}_{1}\bm{V}_{1}^{\mathsf{T}}$ to be the SVD of the optimal rank $k$ approximation and let $\bm{M}_{2}=\bm{B}-\bm{B}_{1}=\bm{U}_{2}\bm{\Sigma}_{2}\bm{V}_{2}^{\mathsf{T}}$ be the remainder. Define $\bm{\Omega}_{1}=\bm{V}_{1}^{\mathsf{T}}\bm{\Omega}$ and $\bm{\Omega}_{2}=\bm{V}_{2}^{\mathsf{T}}\bm{\Omega}$, which are independent Gaussians. Now note that
@@
$\|\bm{B}(\bm{I}-\bm{\Omega}\bm{\Omega}_{1}^{\dagger}\bm{V}_{1}^{T})\|_{\mathsf{F}}$ $=\|\bm{B}(\bm{I}-\bm{V}_{1}^{\mathsf{T}}\bm{V}_{1}^{\mathsf{T}})(\bm{I}-\bm{\Omega}\bm{\Omega}_{1}^{\dagger}\bm{V}_{1}^{\mathsf{T}})\|_{\mathsf{F}}$
$=\sqrt{\|\bm{B}(\bm{I}-\bm{V}_{1}\bm{V}_{1}^{\mathsf{T}})\|_{\mathsf{F}}^{2}+\|\bm{B}(\bm{I}-\bm{V}_{1}\bm{V}_{1}^{\mathsf{T}})\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\bm{V}_{1}^{\mathsf{T}}\|_{\mathsf{F}}^{2}}$
@@
By noting that $\bm{\Omega}_{1}$ and $\bm{\Omega}_{2}$ are independent and using $\mathbb{E}\|\bm{\Sigma}_{2}\bm{\Omega}_{2}\|_{\mathsf{F}}^{2}=q\|\bm{\Sigma}_{2}\|_{\mathsf{F}}^{2}$ [36, Proposition 10.1] and $\mathbb{E}\|\bm{\Omega}_{1}^{\dagger}\|_{2}^{2}\leq\frac{e^{2}q}{(q-k)^{2}-1}$ [57, Lemma 3.1] we have
@@ 4.3 Approximating the diagonal blocks
In this section, we demonstrate how to approximate the diagonal blocks at each iteration of Algorithm 1, while satisfying the condition in Equation 4 of Theorem 3.1. The approach follows the same idea as in [41, Section 4], where the authors assumed that the matrix is exactly HSS. Here, we show that this method is robust even when the matrix is only approximately HSS.
@@
$\bm{U}_{1},\ldots,\bm{U}_{2^{L}}\quad$ where $\text{where }\bm{U}_{i}\in\mathbb{R}^{2k\times k}\text{ has orthonormal columns,}$,
$\bm{V}_{1},\ldots,\bm{V}_{2^{L}}\quad$ where $\text{where }\bm{V}_{i}\in\mathbb{R}^{2k\times k}\text{ has orthonormal columns}.$
@@
$\mathbb{E}\Big{[}\|\bm{B}_{ii}-\bm{D}_{i}-\bm{U}_{i}\bm{U}_{i}^{\mathsf{T}}(\bm{B}_{ii}-\bm{D}_{i})\bm{V}_{i}\bm{V}_{i}^{\mathsf{T}}\|_{\mathsf{F}}^{2}\Big{]}$
@@
Furthermore, noting that $\bm{U}_{i}\bm{U}_{i}^{\mathsf{T}}\bm{D}_{i}\bm{V}_{i}\bm{V}_{i}^{\mathsf{T}}=\bm{0}$ we have and by Pythagoras theorem,
$\|\bm{B}_{ii}-\bm{D}_{i}-\bm{U}_{i}\bm{U}_{i}^{\mathsf{T}}(\bm{B}_{ii}-\bm{D}_{i})\bm{V}_{i}\bm{V}_{i}^{\mathsf{T}}\|_{\mathsf{F}}^{2}$
@@
Taking expectations, using the independence of $\bm{G}_{i},\bm{H}_{i},\bm{\Omega}_{i}$, and $\bm{\Psi}_{i}$ and applying [36, Proposition 10.1-10.2] yields the desired result.
@@ 4.4 A matrix-vector query algorithm for HSS approximation
An obstacle to proving an accuracy guarantee for the algorithm of [41] is the reuse of sketches at each level of the hierarchy. In particular, while at level $L$, one can utilize block nullification to sketch the column and row blocks with Gaussian inputs, this is not the case at subsequent levels. Instead, the implicit sketches are with the same initial Gaussian vectors multiplied by successively more approximated telescoping factors recovered in previous levels [41, Remark 4.3]. This introduces problematic dependencies in the sketching matrices across levels.
@@
12:   $\bm{U}^{(\ell)}=\text{blockdiag}(\bm{U}_{1}^{(\ell)},\ldots,\bm{U}_{2^{\ell}}^{(\ell)})$.
13:   $\bm{V}^{(\ell)}=\text{blockdiag}(\bm{V}_{1}^{(\ell)},\ldots,\bm{V}_{2^{\ell}}^{(\ell)})$.
14:   $\bm{D}^{(\ell)}=\text{blockdiag}(\bm{D}_{1}^{(\ell)},\ldots,\bm{D}_{2^{\ell}}^{(\ell)})$.
15:   $\bm{A}^{(\ell)}=(\bm{U}^{(\ell)})^{\mathsf{T}}\left(\bm{A}^{(\ell+1)}-\bm{D}^{(\ell)}\right)\bm{V}^{(\ell)}$.
@@
First, note that by the discussion in Subsection 4.2, $\bm{Y}_{i}^{(\ell)}\bm{P}_{i}^{(\ell)}$ is a sketch of $r_{i}(\bm{A}^{(\ell+1)})$ with a standard Gaussian random matrix with $s-2k$ columns. Hence, by Theorem 4.2 with $q=s-2k$ we have
@@
Note that the matrices $\bm{A}^{(\ell+1)}$ for $\ell=L,\ldots,1$ are never formed explicitly, but are only accessed through matvecs with $\bm{A}$ as described in Eq. 15. Hence, the cost of computing matvecs with $\bm{A}^{(\ell+1)}$ comes from performing the operations in the recursion Eq. 15 and the matvecs with $\bm{A}$. Let $c_{\ell+1}$ be the cost of computing $s$ matvecs with $\bm{A}^{(\ell+1)}$ and its transpose using the recursion in (15), and let $C$ be the cost of computing $s$ matvecs with $\bm{A}$ and its transpose. Then, we have $c_{\ell+1}=O(2^{\ell}k^{2}s)+c_{\ell+2}$ with $c_{L+1}=C$. Consequently, summing over $\ell=1,\ldots,L$ and using $L=\log_{2}(N/k)$ the total cost arising from the computing matrix-vector products with $\bm{A}^{(\ell+1)}$ and its transposes is $O(NksL+LC)$. For a fixed level $\ell$, the additional runtime required by the inner for-loop and sampling the random Gaussian matrices is $O(2^{\ell}(k^{2}s+ks^{2}))$. Summing over $\ell$ and using $L=\log_{2}(N/k)$ yields an additional runtime of $O(Nks+Ns^{2})=O(Ns^{2})$, assuming $s\geq k$. Hence, the total cost of Algorithm 2 is $O(sL)$ matvecs with $\bm{A}$ and its transpose, and $O(NksL+Ns^{2})$ additional runtime.
@@ 5 Numerical experiments
In this section, we compare our Algorithm 2 to [41, Algorithm 3.1], on which it is based. All experiments were performed on a Macbook, and Python scripts to reproduce the figures in this section can be found at https://github.com/NoahAmsel/HSS-approximation.
Recall that [41, Algorithm 3.1] reuses the same random sketching matrix across levels. Therefore, it only requires only $O(k)$ matvec queries with $\bm{A}$, which compares favorably to $O(Lk)$ matvec queries required by Algorithm 2. However, as discussed in Subsection 4.4, the reuse of randomness leads to difficult to analyze dependencies in error terms across levels. As such, no strong theoretical approximation guarantees are known for [41, Algorithm 3.1] and, as we will see shortly, for a fixed sketch size, the algorithm is indeed less accurate than our variant that uses fresh sketches for each level. It thus remains an interesting open question to develop an $O(k)$ query algorithm that matches the theoretical performance of Algorithm 2.
@@
A matrix $\bm{M}$ is said to have bandwidth $b$ if $\bm{M}_{ij}=0$ whenever $|i-j|>b$. Let $N=2^{L+1}$ and let $\bm{M}\in\mathbb{R}^{N\times N}$ have bandwidth $b=2k+1$. Then $\bm{A}:=\bm{M}^{-1}$ is HSS with rank $2k$ [34, Section 3.9]. We generate random symmetric banded $\bm{M}$ and approximate $\bm{A}$ with a matrix $\bm{B}\in\operatorname{HSS}(L,k)$ (i.e., smaller rank than needed to achieve zero error). Matrix-vector products with $\bm{A}$ are performed using a sparse Cholesky factorization of $\bm{M}$. In our experiments, we set $b=17,N=4096,L=9$, and $k=8$.
@@
[…]
@@
Figure 3: Experimental results of HSS approximation on four model problems. Each line shows a different algorithm: Refer to captionAlgorithm 2 (fresh sketches at each level), Refer to caption [41, Algorithm 3.1] (reused sketches across levels; basis found by pivoted QR), Refer to captionreused sketches with the basis found by SVD, Refer to captionthe version of Algorithm 1 for explicit matrices (page 3). Shaded regions show range over ten runs, though it is sometimes imperceptible.
@@ 5.1 Results
We implement and run both [41, Algorithm 3.1] and Algorithm 2 on the test problems described above. To isolate the effect of the reuse of random sketches, we also address another implementation difference between the algorithms. Specifically, [41] computes the bases $\bm{U}^{(\ell)}$ and $\bm{V}^{(\ell)}$ using a pivoted QR decomposition, while Algorithm 2 uses the SVD. To remove any influence this minor difference in implementation may have on the results, we also implement a modified version of [41, Algorithm 3.1] in which we replace the pivoted QR step with an SVD. We run Algorithm 2 (fresh sketches), [41, Algorithm 3.1] (reused sketches, QR), and our modification of the latter (reused sketches, SVD) with a range of sketch sizes $s$. Following Theorem 4.1, we set the sketch size $s$ to be at least $23k+$ 2. For each $s$, we run each algorithm ten times and measure the relative errors $\|\bm{A}-\bm{B}\|_{\mathsf{F}}/\|\bm{A}\|_{\mathsf{F}}$.
@@
Figure 3 plots the results of all experiments. For each test matrix, the left side plots the relative approximation error of each algorithm as a function of the sketch size $s$. The right side plots it as a function of the total number of matrix-vector queries with $\bm{A}$ and $\bm{A}^{\mathsf{T}}$ across all $L$ levels of the algorithm (“Total Queries”), which is $4s$ for [41, Algorithm 3.1] and $4Ls$ for Algorithm 2. Lines show the mean over ten runs of each algorithm, and shaded regions show the range.
The results show that the reuse of randomness in [41, Algorithm 3.1] does have an adverse effect on the error as compared to Algorithm 2. However, for a fixed number of total matrix-vector products, [41, Algorithm 3.1] has lower error than Algorithm 2. Hence, while necessary for our theoretical guarantees, for the problems tested, reusing sketches per level does not lead to a more efficient algorithm. This finding leaves us hopefully that it might be possible to improve on our matvec bound of $O(N\log(N/k))$ by the logarithmic factor, possibly via a more refined analysis of sketch reuse.

### v2

@@ 4.1 Block nullification
Since $\bm{\Omega}_{i}^{(\ell)}\in\mathbb{R}^{2k\times s}$, it has a nullspace of dimension at least s2k>ks-2k>k. Therefore, we can obtain a matrix $\bm{P}_{i}^{(\ell)}$ whose columns form an orthonormal basis for the nullspace of $\bm{\Omega}_{i}^{(\ell)}$. Then, $\bm{\Omega}_{i}^{(\ell)}\bm{P}_{i}^{(\ell)}=\bm{0}$. Moreover, for all $j\neq i$, the unitary invariance of Gaussian matrices together with the mutual independence of $\bm{\Omega}_{i}$ and $\bm{\Omega}_{j}$ imply that $\bm{\Omega}_{j}^{(\ell)}\bm{P}_{i}^{(\ell)}$ are independent Gaussian matrices. Therefore, if we define
@@ 4.2 Low-rank approximation via projection-cost-preserving sketches
Define $\bm{B}_{1}=\bm{U}_{1}\bm{\Sigma}_{1}\bm{V}_{1}^{\mathsf{T}}$ to be the SVD of the optimal rank-$k$ approximation and let $\bm{M}_{2}=\bm{B}-\bm{B}_{1}=\bm{U}_{2}\bm{\Sigma}_{2}\bm{V}_{2}^{\mathsf{T}}$ be the remainder. Define $\bm{\Omega}_{1}=\bm{V}_{1}^{\mathsf{T}}\bm{\Omega}$ and $\bm{\Omega}_{2}=\bm{V}_{2}^{\mathsf{T}}\bm{\Omega}$, which are independent Gaussians. Now note that [^7] This proof technique is related to the analysis of the randomized SVD and generalized Nyström approximation in [59, Section 3.3-3.4], which was inspired by an observation in [66, Section 4]; the latter attributes the observation to Ilse Ipsen.
@@
$\|\bm{B}(\bm{I}-\bm{\Omega}\bm{\Omega}_{1}^{\dagger}\bm{V}_{1}^{T})\|_{\mathsf{F}}$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\|\bm{B}(\bm{I}-\bm{V}_{1}\bm{V}_{1}^{\mathsf{T}})(\bm{I}-\bm{\Omega}\bm{\Omega}_{1}^{\dagger}\bm{V}_{1}^{\mathsf{T}})\|_{\mathsf{F}}}$
$=\sqrt{\|\bm{B}(\bm{I}-\bm{V}_{1}\bm{V}_{1}^{\mathsf{T}})\|_{\mathsf{F}}^{2}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\|\bm{B}(\bm{I}-\bm{V}_{1}\bm{V}_{1}^{\mathsf{T}})\bm{\Omega}\bm{\Omega}_{1}^{\dagger}\bm{V}_{1}^{\mathsf{T}}\|_{\mathsf{F}}^{2}}}$
@@
By noting that $\bm{\Omega}_{1}$ and $\bm{\Omega}_{2}$ are independent and using $\mathbb{E}\|\bm{\Sigma}_{2}\bm{\Omega}_{2}\|_{\mathsf{F}}^{2}=q\|\bm{\Sigma}_{2}\|_{\mathsf{F}}^{2}$ [38, Proposition 10.1] and $\mathbb{E}\|\bm{\Omega}_{1}^{\dagger}\|_{2}^{2}\leq\frac{e^{2}q}{(q-k)^{2}-1}$ [59, Lemma 3.1] we have
@@ 4.3 Approximating the diagonal blocks
In this section, we demonstrate how to approximate the diagonal blocks at each iteration of Algorithm 1, while satisfying the condition in Equation 4 of Theorem 3.1. The approach follows the same idea as in [43, Section 4], where the authors assumed that the matrix is exactly HSS. Here, we show that this method is robust even when the matrix is only approximately HSS.
@@
$\bm{U}_{1},\ldots,\bm{U}_{2^{L}}\quad$ $\text{where }\bm{U}_{i}\in\mathbb{R}^{2k\times k}\text{ has orthonormal columns,}$
$\bm{V}_{1},\ldots,\bm{V}_{2^{L}}\quad$ $\text{where }\bm{V}_{i}\in\mathbb{R}^{2k\times k}\text{ has orthonormal columns}.$
@@
$\hskip-40.00006pt\mathbb{E}\Big{[}\|\bm{B}_{ii}-\bm{D}_{i}-\bm{U}_{i}\bm{U}_{i}^{\mathsf{T}}(\bm{B}_{ii}-\bm{D}_{i})\bm{V}_{i}\bm{V}_{i}^{\mathsf{T}}\|_{\mathsf{F}}^{2}\Big{]}$
@@
Furthermore, noting that $\bm{U}_{i}\bm{U}_{i}^{\mathsf{T}}\bm{D}_{i}\bm{V}_{i}\bm{V}_{i}^{\mathsf{T}}=\bm{0}$ and applying Pythagoras theorem yields
$\hskip-40.00006pt\|\bm{B}_{ii}-\bm{D}_{i}-\bm{U}_{i}\bm{U}_{i}^{\mathsf{T}}(\bm{B}_{ii}-\bm{D}_{i})\bm{V}_{i}\bm{V}_{i}^{\mathsf{T}}\|_{\mathsf{F}}^{2}$
@@
Taking expectations, using the independence of $\bm{G}_{i},\bm{H}_{i},\bm{\Omega}_{i}$, and $\bm{\Psi}_{i}$ and applying [38, Proposition 10.1-10.2] yields the desired result.
@@ 4.4 A matrix-vector query algorithm for HSS approximation
An obstacle to proving an accuracy guarantee for the algorithm of [43] is the reuse of sketches at each level of the hierarchy. In particular, while at level $L$, one can utilize block nullification to sketch the column and row blocks with Gaussian inputs, this is not the case at subsequent levels. Instead, the implicit sketches are with the same initial Gaussian vectors multiplied by successively more approximated telescoping factors recovered in previous levels [43, Remark 4.3]. This introduces problematic dependencies in the sketching matrices across levels.
@@
[…]
@@
First, note that by the discussion in Subsection 4.2, $\bm{Y}_{i}^{(\ell)}\bm{P}_{i}^{(\ell)}$ is a sketch of $r_{i}(\bm{A}^{(\ell+1)})$ with a standard Gaussian random matrix with s2ks-2k columns. Hence, by Theorem 4.2 with q=s2kq=s-2k we have
@@
Note that the matrices $\bm{A}^{(\ell+1)}$ for $\ell=L,\ldots,1$ are never formed explicitly, but are only accessed through matvecs with $\bm{A}$ as described in Eq. 15. Hence, the cost of computing matvecs with $\bm{A}^{(\ell+1)}$ comes from performing the operations in the recursion Eq. 15 and the matvecs with $\bm{A}$. Let $c_{\ell+1}$ be the cost of computing $s$ matvecs with $\bm{A}^{(\ell+1)}$ and its transpose using the recursion in (15), and let $C$ be the cost of computing $s$ matvecs with $\bm{A}$ and its transpose. Then, we have $c_{\ell+1}=O(2^{\ell}k^{2}s)+c_{\ell+2}$ with $c_{L+1}=C$. Consequently, summing over $\ell=1,\ldots,L$ and using $L=\log_{2}(N/k)$ the total cost of computing these products is $O(NksL+LC)$. For a fixed level $\ell$, the additional runtime required by the inner for-loop and sampling the random Gaussian matrices is $O(2^{\ell}(k^{2}s+ks^{2}))$. Summing over $\ell$ and using $L=\log_{2}(N/k)$ yields an additional runtime of $O(Nks+Ns^{2})=O(Ns^{2})$, assuming $s\geq k$. Hence, the total cost of Algorithm 2 is $O(sL)$ matvecs with $\bm{A}$ and its transpose, and $O(NksL+Ns^{2})$ additional runtime.
@@ 5 Numerical experiments
In this section, we compare our Algorithm 2 to [43, Algorithm 3.1], on which it is based. All experiments were performed on a Macbook, and Python scripts to reproduce the figures in this section can be found at https://github.com/NoahAmsel/HSS-approximation.
Recall that [43, Algorithm 3.1] reuses the same random sketching matrix across levels. Therefore, it only requires only $O(k)$ matvec queries with $\bm{A}$, which compares favorably to $O(Lk)$ matvec queries required by Algorithm 2. However, as discussed in Subsection 4.4, the reuse of randomness leads to difficult to analyze dependencies in error terms across levels. As such, no strong theoretical approximation guarantees are known for [43, Algorithm 3.1] and, as we will see shortly, for a fixed sketch size, the algorithm is indeed less accurate than our variant that uses fresh sketches for each level. It thus remains an interesting open question to develop an $O(k)$ query algorithm that matches the theoretical performance of Algorithm 2.
@@
A matrix $\bm{M}$ is said to have bandwidth $b$ if $\bm{M}_{ij}=0$ whenever |ij|>b|i-j|>b. Let $N=2^{L+1}$ and let $\bm{M}\in\mathbb{R}^{N\times N}$ have bandwidth $b=2k+1$. Then $\bm{A}:=\bm{M}^{-1}$ is HSS with rank $2k$ [36, Section 3.9]. We generate random symmetric banded $\bm{M}$ and approximate $\bm{A}$ with a matrix $\bm{B}\in\operatorname{HSS}(L,k)$ (i.e., smaller rank than needed to achieve zero error). Matrix-vector products with $\bm{A}$ are performed using a sparse Cholesky factorization of $\bm{M}$. In our experiments, we set $b=17,N=4096,L=9$, and $k=8$.
@@
(22) $\frac{1}{2}\sigma(x)-\frac{1}{2\pi}\int_{\partial D}\frac{n(x)\cdot(x-y)}{\|x-y\|^{2}}\sigma(y)dy=f(x)\qquad\text{for all }x\in\partial D.$
@@
Figure 3: Experimental results of HSS approximation on four model problems. Each line shows a different algorithm: Refer to captionAlgorithm 2 (fresh sketches at each level; requires $4sL$ matvec queries), Refer to caption [43, Algorithm 3.1] (reused sketches across levels; basis found by pivoted QR; requires $4s$ matvec queries), Refer to captionreused sketches with the basis found by SVD (requires $4s$ matvec queries), Refer to captionthe version of Algorithm 1 for explicit matrices (page 3). Shaded regions show range over ten runs, though it is sometimes imperceptible.
@@ 5.1 Results
We implement and run both [43, Algorithm 3.1] and Algorithm 2 on the test problems described above. To isolate the effect of the reuse of random sketches, we also address another implementation difference between the algorithms. Specifically, [43] computes the bases $\bm{U}^{(\ell)}$ and $\bm{V}^{(\ell)}$ using a pivoted QR decomposition, while Algorithm 2 uses the SVD. To remove any influence this minor difference in implementation may have on the results, we also implement a modified version of [43, Algorithm 3.1] in which we replace the pivoted QR step with an SVD. We run Algorithm 2 (fresh sketches), [43, Algorithm 3.1] (reused sketches, QR), and our modification of the latter (reused sketches, SVD) with a range of sketch sizes $s$. Following Theorem 4.1, we set the sketch size $s$ to be at least $3k+2$. For each $s$, we run each algorithm ten times and measure the relative errors $\|\bm{A}-\bm{B}\|_{\mathsf{F}}/\|\bm{A}\|_{\mathsf{F}}$.
@@
Figure 3 plots the results of all experiments. For each test matrix, the left side plots the relative approximation error of each algorithm as a function of the sketch size $s$. The right side plots it as a function of the total number of matrix-vector queries with $\bm{A}$ and $\bm{A}^{\mathsf{T}}$ across all $L$ levels of the algorithm (“Total Queries”), which is $4s$ for [43, Algorithm 3.1] and $4Ls$ for Algorithm 2. Lines show the mean over ten runs of each algorithm, and shaded regions show the range.
The results show that the reuse of randomness in [43, Algorithm 3.1] does have an adverse effect on the error as compared to Algorithm 2. However, for a fixed number of total matrix-vector products, [43, Algorithm 3.1] has lower error than Algorithm 2. Hence, while necessary for our theoretical guarantees, for the problems tested, using independent sketches per level does not lead to a more efficient algorithm. In practice, we thus recommend reusing sketches across levels. This finding leaves us hopeful that it might be possible to improve on our matvec bound of $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}O(k\log(N/k))$} by the logarithmic factor via a more refined analysis of sketch reuse. Furthermore, the experiments indicate that the modification of [43, Algorithm 3.1], which computes the bases using the SVD, performs better than the version using the pivoted QR algorithm.
@@
Acknowledgments
We thank the referees for helpful comments on this work. We are indebted to an anonymous referee for outlining how to prove the existence of an optimal HSS approximation.